OpenStudy (anonymous):

Calculate the slope between the points (-6, 9) and (4, 11).

OpenStudy (anonymous):

0.2, i'll draw

hero (hero):

I like to use y = mx + b to solve these. All you do is simply enter the points to the formula and perform the following steps to solve for m: 9 = -6m + b 1 = 4m + b Isolate b: b = 9 + 6m b = 1 - 4m Set b = b 9 + 6m = 1 - 4m 6m + 4m = 1 - 9 10m = -8 10/-8 = m -5/4 = m

hero (hero):

@mickey4691, be advised that OS prefers when users posts steps, not answers alone.